Bill, an investor, purchases a six-month (182-day) T-bill with a $11,000 par value for $10,400. If Moe holds the T-bill to maturity what is his bond equivalent yield?
Answer #1

We see that the Bond Equivalent yield is given as equal to=(Face Value/Purchase Price-1)*365/d=(11000/10400-1)*365/182=11.5702%